twice the difference of a number and 7 is equal to three times the sum of the number and 6 find the number
Found 2 solutions by robertb, stanbon:
Answer by robertb(5830)

2(x-7) = 3(x +6),
2x-14 = 3x+18,
-x = 32,
or x = -32.

Answer by stanbon(75887)

twice the difference of a number and 7 is equal to three times the sum of the number and 6 find the number..
Equation:
2(n-7) = 3(n+6)
2n-14 = 3n+18
n = -32
--------------
Cheers,
Stan H.